OpenStudy (anonymous):

determine the three cubic roots of 8.give the answer in the form of a+bj.also give the main root in the exponential form

OpenStudy (callisto):

Is this what's stated in your question? \[\sqrt[3]{8} ?\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

the other two roots i suppose would be the imaginary ones :/

OpenStudy (callisto):

\[\sqrt[3]{8} = (2^3)^{\frac{1}{3}} =?\]

OpenStudy (anonymous):

yes other two imaginery

OpenStudy (lgbasallote):

you didnt happen to copy-paste that from http://answers.yahoo.com/question/index?qid=20071207173154AACVdWs did you @HairtUB ?

OpenStudy (anonymous):

Aaah busted! haha but I swear that's the only time I copied a work! Damn, I was thinking of sourcing it but didn't think anyone would notice...oh well I guess that's what you get for plagiarizing...I learnt my lesson.

OpenStudy (anonymous):

Alright this time I'll source: I got an answer from http://answers.yahoo.com/question/index?qid=20071207173154AACVdWs Let's solve the equation x³ = 8. We have x³-8 = 0 (x-2)(x²+2x + 4) = 0 So the real root is x = 2 and the complex cube roots are x = -1 + i√3 and x = -1 -i√3.

OpenStudy (unklerhaukus):